# Scalable Ab Initio Electronic Structure Methods with Near Chemical Accuracy for Main Group Chemistry

### v1

@@ Abstract
This study evaluates the precision of widely recognized quantum chemical methodologies, CCSD(T), DLPNO-CCSD(T) and localized ph-AFQMC, for determining the thermochemistry of main group elements. DLPNO-CCSD(T) and localized ph-AFQMC, which offer greater scalability compared to canonical CCSD(T), have emerged over the last decade as pivotal in producing precise benchmark chemical data. Our investigation includes closed-shell, neutral molecules, focusing on their heat of formation and atomization energy sourced from four specific small molecule datasets. Firstly, we selected molecules from the G2 1 and G3 2 datasets, noted for their reliable experimental heat of formation data. Additionally, we incorporate molecules from the W4-11 3 and W4-17 4 sets, which provide high-level theoretical reference values for atomization energy at 0 K. Our findings reveal that both DLPNO-CCSD(T) and ph-AFQMC methods are capable of achieving a root-mean-square deviation (RMSD) of less than 1 kcal/mol across the combined dataset, aligning with the threshold for chemical accuracy. Moreover, we make efforts to confine the maximum deviations within 2 kcal/mol, a degree of precision that significantly broadens the applicability of these methods in fields such as biology and materials science.
@@
1 I. Introduction
Quantum chemical methods have seen significant improvements in accuracy and computational efficiency when applied to the chemistry of main group elements over the last three decades. Density functional methods, which scale formally with system size $N$ as $N^{4}$ or $N^{3}$ and in practice as $N^{2}$ or even $N$ (for very large systems), can routinely be applied to systems with hundreds to thousands of atoms, with the best functionals providing an average unsigned error on the order of 2–3 kcal/mol for atomization energies of small molecules.5 Alternatively, coupled cluster with perturbative triples (CCSD(T)) has been integrated into numerous quantum chemistry software packages in a robust manner, offering average error rates close to 1 kcal/mol for atomization energies of the same category of small molecules; however, the computational cost scales as $N^{7}$.6, 7 Benchmark methodologies that incorporate higher-order coupled cluster terms and more elaborate treatment of core-valence interactions are capable of producing results reliably within 1 kcal/mol deviation from experimental values. 8 However, the high computational demand of these approaches limits their application to very small molecular systems.
While the accuracy that can be achieved with modern DFT approaches is extremely impressive, DFT calculations on large data sets reveal a significant number of outliers with errors significantly larger than the 2–3 kcal/mol cited above as the average unsigned error. Importantly, outliers can be obscured when large data sets are used and only the average errors are reported.5 In a subsequent paper, we will examine in detail the outlier distribution obtained for a range of modern functionals when compared to benchmark methods and curated experimental results. For the present purposes it is sufficient to note that more work remains to be done to improve the robustness of DFT approaches across a wide range of chemistry, even for main group molecules. Moreover, systems containing transition metals can be prone to a higher incidence of outliers.9
@@
However, in order for wavefunction-based methods to effectively address complex systems, an approach is needed which scales better with system size than the $N^{7}$ of conventional CCSD(T). The past decade has seen the development of two notable methods that address this need, both leveraging the concept of orbital localization—a technique tracing back to Pulay’s work in the 1980s.10 The first is localized coupled cluster (e.g. L-CCSD(T)), the most widely used implementation of which is the DLPNO algorithm of Neese and coworkers.11, 12, 13 The formal scaling of DLPNO-CCSD(T) is $N^{3}$ 14 and extremely impressive timing and accuracy numbers over a wide range of systems (and particularly those restricted to main group chemistry) have been published in the past 5 years.15, 16, 17, 18, 19, 20 DLPNO and related methods represent a revolution in quantum chemical technology as it is now possible to obtain something close to CCSD(T) quality results for systems containing on the order of 100 atoms.
@@
The second approach is auxiliary field quantum Monte Carlo (AFQMC). The AFQMC algorithm was originally developed in the physics community, but it is only in the past 5 years that significant progress has been made in creating a scalable version of the methodology for the ab initio study of molecules. There are several different implementations currently in use. 21, 22, 23, 24. In the present paper, we will focus on two of these implementations. The first is a GPU implementation developed in our groups 25 that exploits localized orbitals 26 in a fashion similar to that employed in L-CCSD(T), reducing the formal scaling of AFQMC from $N^{4}$ to $N^{3}$. We refer to this approach as L-AFQMC. The second 27, 28 is a CPU-based code optimized to enable systematic convergence of the bias in AFQMC energies to near-exact accuracy by using a large number of determinants in the trial function. We designate this implementation as W-AFQMC since it is based on use of the generalized Wick’s theorem. As yet, this method has not been formulated in a localized representation, although work in that direction is ongoing.29 To avoid high computational costs, here we employ W-AFQMC to resolve discrepancies identified in DLPNO-CCSD(T) and L-AFQMC data.
@@
While the computational cost scaling of L-AFQMC and L-CCSD(T) with system size is similar, L-CCSD(T) is considerably faster due to a smaller prefactor. The advantage of AFQMC is that it is formally exact in the limit of the exact trial wavefunction, and in practice, multireference electronic states can often be readily converged due to the ease of utilizing a multiconfigurational trial function.30, 31, 32, 33, 28 This is more crucial for transition metal containing systems than for main group molecules, but there are still main group cases where AFQMC can achieve demonstrably greater accuracy with scalable trial wavefunctions.34, 35 While such trials allow systematic convergence of the bias in phaseless AFQMC (ph-AFQMC), this accuracy comes at the cost of greater computational expense. Designing protocols for generating trials that strike a desired balance between accuracy and cost is an active area of research. Based on our testing with the benchmark sets, we present two approaches with different cost-accuracy trade-offs.
@@
We believe that having two scalable benchmark methodologies with distinct theoretical frameworks offers substantial benefits. These benefits extend not only to the generation of data for evaluating and parameterizing DFT functionals but also to their direct application to challenging systems, like the manganese cluster in Photosystem II. As an example, in one of our recent papers we investigated reactions of organolithium systems relevant to lithium ion batteries.36 We performed DLPNO-CCSD(T) and L-AFQMC calculations to look at both reaction energies and barrier heights. The agreement between the two approaches was remarkably good (a few kcal/mol) across the various reactions that we investigated. We were therefore able to settle on benchmark numbers and use those to evaluate many different DFT functionals, discovering that only a few were able to reproduce the benchmark results reliably. We were then able to use the preferred functionals in computing energies for a large set of organolithium cluster geometries, which we then utilized in parametrizing a machine learning force field (MLFF) for carrying out simulations of lithium ion battery electrolytes.37 While there was no reason, a priori, to doubt the performance of DLPNO-CCSD(T) for these systems, there is very little experimental data available for comparisons, and the validation by a second independent benchmark approach provided a much higher degree of confidence in the results than would otherwise have been possible.
@@
In the present paper, our goal is to provide an assessment of the accuracy of both DLPNO-CCSD(T) and L-AFQMC for main group chemistry atomization energies. We have chosen to focus on atomization energies because (a) a relatively large and reliable data set of benchmark experimental and theoretical values is available for a range of small molecules and (b) atomization energies are one of the most difficult properties for electronic structure methods to compute to high precision, due to the large changes in correlation energy upon atomization, and the relatively minimal cancellation of error. We combine four data sets: the G21, 38 and G32, 39 sets of Pople and coworkers, which claim to have experimental atomization energies that are accurate to 1 kcal/mol or better, and the W4-11 3 and W4-17 4 (TAE - total atomization energy) data sets of Karton, Martin and coworkers, which employ very high level (and hence expensive) theoretical methods to achieve the same level of reliability. The W4-17 set, the latest iteration of the W4 sets, is an extension of the W4-11 set, as is G3 an extension of G2 with the addition of larger molecules. W4-17, however, is restricted to molecules with no more than 8 heavy atoms. On the other hand, the G3 data set contains larger molecules than the W4-17 heavy atom limit and therefore provides a test of how the quantum chemical methods perform for larger systems. There are (as far as we have been able to ascertain) no cases where the errors in the W4-11 and W4-17 results exceed the proposed error bars. In contrast, we have had to update a number of the G2 and G3 experimental values with more recent benchmark values.
@@
The paper is organized as follows. In Section II, we discuss the data sets including our updating of a number of the reference values in the G2 and G3 sets. Section III discusses computational methods including our scalable formulation of AFQMC, DLPNO-CCSD(T), and CCSD(T) methodology (enabling comparisons for the larger systems), basis sets, core-valence corrections, scalar relativistic corrections, extrapolation to the complete basis set limit, and treatment of atomic energies. Section IV presents the results for all three methods as compared to the relevant experimental or theoretical benchmarks and discusses their implications. In Section V, we conclude by summarizing our results and discussing future directions. In general, the performance of both DLPNO-CCSD(T) and L-AFQMC are quite robust, with only a few apparent outliers above our targeted 2 kcal/mol threshold. We investigate these outliers using more accurate trials with the W-AFQMC method. This enables us to identify outliers arising from the approximations in the methodology, as opposed to cases that are most likely errors or uncertainties in a few of the experimental reference data. Using this targeted convergence of ph-AFQMC, we are able to produce high-quality atomization energies while minimizing cost over a large dataset.
2 II. Datasets
@@ 2.1 Reference Values
Table 1: New vs old reference values for heat of formation. The reference values in the column ‘Previous $\Delta_{f}H{(298\,\text{K})}$’ are the same as used in the original G2/G3 test set papers.38, 39 All values are reported in kcal/mol. Where the source reports an experiment uncertainty, we have included the uncertainty in the table along with the value.
Reference values for the W4 sets are high-level theoretical atomization energies at 0 K, excluding zero point energy (ZPE) (labeled as the property TAEe by Karton et al. 4). The reference values for the G2 and G3 sets are experimental heat of formation 298 K, and the same as those used by Curtiss et al. 1, 2. However, there are some exceptions where we have found conflicting values, as summarized by Table 1 where each of the sources of the updated reference are listed. The reported reference values from ATcT40, 41 postdate the G2 and G3 papers and collates the most recent experiments and theory from various sources42. Therefore, where ATcT data is available and conflicting with the reference used by Curtiss et al., we instead use the ATcT value. In addition, for the cases of AlCl3 and AlF3 we found disagreement between the experimental heats of formation and the W4 reported atomization energy. Moreover, the heat of formation values of these molecules are not present in the ATcT database. Thus, for the updated reference values of these molecules, we have converted the W4 atomization values into heats of formation at 298 K using temperature corrections from DFT (refer to Section III for details). For the three cases LiF, azulene, and benzoquinone, where no ATcT or W4 value is available, instead of the reference reported by Curtiss, we use those reported by the NIST database CCCBDB43. As shown in Table 1, the two sources also give conflicting heats of formation for these three molecules. All of our benchmark wavefunction methods yield results that are within 2 kcal/mol of the latter source, rather than the former reference values. Although this choice is not based on information about the reference alone, CCSD(T) and AFQMC offer independent evaluations of the experimental data, especially in the case of conflict. In particular, we converge AFQMC with respect to the number of determinants for these cases using W-AFQMC. We refer the reader to Section III for the method of obtaining deviations against reference values.
3 III. Computational Details
3.1 Phaseless AFQMC Formulation
@@
$h_{pq}$ are one-electron integrals and $V_{pqrs}=\left(pq|rs\right)=\langle pq|rs\rangle{}{}$ are two-electron integrals. Numerically, we propagate
@@
where $p(\mathbf{x})$ is a Gaussian probability density function and $\hat{B}(\mathbf{x})$ is a one-body propagator depending on the auxiliary fields $\mathbf{x}$. This multidimensional integral is evaluated using Monte Carlo importance sampling to obtain a stochastic representation of the wave function. For a more in-depth description of AFQMC, we refer the reader to these review articles 44, 45.
@@
Due to the fermionic sign problem, the signal-to-noise ratio generally decays exponentially during the imaginary time propagation. It is possible to eliminate the sign problem using a constraint at the expense of a bias in the resulting energies. In this work, we use a constraint referred to as the phaseless approximation (ph-AFQMC), where the phase of the walkers is restricted according to a trial wavefunction. The bias induced by the trial wavefunction can be systematically reduced by improving its quality, for example, by increasing the active space or number of determinants included in the trial. The bias is formally zero in the limit of the exact trial (see the next subsection, Section IV, and SI Section 3 for trial details). We refer the reader to our previous work26 for our approach to localized ph-AFQMC (L-AFQMC) which involves compressing the electron repulsion integrals in the localized orbital basis. Effectively, the scaling for the energy evaluation, the steepest scaling step with system size, is $N^{2}M+N_{\text{det}}N^{2}$ (with a prefactor depending on the compressed electron repulsion integral tensor), where $M$ is the number of basis functions, $N$ is the number of electrons, and $N_{\text{det}}$ is the number of determinants. See SI Section 4 for an estimate of localization error. For the practical deployment of L-AFQMC, we have developed two protocols, AFQMC 0 and AFQMC I, which are discussed in detail below as well as in Section IV. AFQMC I is a scalable AFQMC protocol (scaling $\sim N^{3}$) that achieves an accuracy comparable to that of DLPNO-CCSD(T), albeit with a significantly larger prefactor for the computation time. AFQMC 0 uses a black-box but less elaborate trial function, but is less accurate, particularly for molecules with significant multireference character.
@@
We also present results for a selected subset of molecules computed with another implementation of AFQMC 27, 28. This method, which uses a generalized Wick’s theorem approach to efficiently evaluate energies with mutlideterminantal trials, will be referred to in what follows as W-AFQMC. The advantage of this approach is that it scales as $MN_{\text{det}}+N^{2}M^{2}$, which allows the use of a larger number of determinants in the trial wavefunction at an accessible computational cost. It enables one to converge the phaseless bias to the near-exact limit in a given basis set by increasing the number of active space orbitals and determinants included in trial wavefunctions. We use W-AFQMC to calculate energies of the outliers obtained from AFQMC I using on the order of 10,000 determinants and refer to the results in which the L-AFQMC outliers are replaced by the W-AFQMC results as AFQMC II (along with a select few other cases, see SI Section 3). This approach helps us to more confidently address the question, discussed in detail below, as to whether the AFQMC I outliers are due to the phaseless bias or more likely the result of errors in the reference data. We also run apparent DLPNO-CCSD(T) and CCSD(T) outliers with AFQMC II to assess their status.
3.2 AFQMC Trial Generation
@@
In this work, we use a procedure to generate multi-determinant HCI (heat bath configuration interaction) and HCISCF trials for the entire dataset. The CAS family of trials (in this work, we use the HCI solver46, 47) provides a robust way of including static correlation in the reference of AFQMC and has been shown21, 30 to generally perform more accurately than single determinant trials. Nonetheless, the selection of active space for these methods is non-trivial. Akin to multireference perturbation theory methods, a common practice is to use an active space, usually minimal, based on chemical intuition and to pick the leading determinants from the expansion in this active space. However, this does not systematically provide chemically accurate energies. Here, we generally follow a two-step process to select the active space in a relatively automated way that can be applied to large datasets. First, HCI is performed on a “valence” active space, selected based on the atomic composition of the molecule. See SI Section 3 and main text Section IV for the considerations for selecting this active space. Using the spatial 1-RDM of the resulting state, we calculate the natural orbitals and their occupation numbers (NOONs) in the natural orbital basis. We choose a subset of active orbitals from this set based on a NOON threshold ($\delta$) as $\delta\leq\text{NOON}\leq 2-\delta$. This procedure is often used to flag the more correlated group of orbitals in quantum calculations. A second HCI calculation (or HCISCF for W-AFQMC trials, refer to the SI Section 3) is then performed with the second smaller active space, and this forms the final trial wavefunction. Generally, we choose the number of determinants necessary to retain 99.5% of the CI weight from this final trial wavefunction unless indicated otherwise. Briefly, AFQMC 0 is fully automated and uses L-AFQMC, and thresholds are chosen to be loose, and for the outliers with AFQMC 0, AFQMC I combines trials with larger active spaces, and AFQMC II in turn combines trials with even larger active spaces run with W-AFQMC. This progression gives some indication of the number of determinants required to converge each molecule, but not much more than necessary. For more details, refer to the discussion in Section IV and SI Section 3.
@@
All L-AFQMC trials were generated with the PySCF package48 where we obtain the Hamiltonian and electron repulsion integrals. The HCI trials are obtained using Dice46, 47, 49 in conjunction with PYSCF. L-AFQMC energy is measured in blocks of size 0.1 Ha-1 of 20 timesteps each (each timestep being 0.005 Ha-1). In total, we propagate for between 2000 and 3000 such blocks for each molecule, with 1920 total walkers. W-AFQMC calculations used the same time-step of 0.005 Ha-1 for molecules containing only first row atoms, and used a time step of 0.0025 Ha-1 for those with heavier atoms. For W-AFQMC, we use 250 walkers and propagate for 1000 blocks of 50 timesteps each.
3.3 CCSD(T) and DLPNO-CCSD(T)

### v3

@@ Abstract
This study evaluates the precision of widely recognized quantum chemical methodologies, CCSD(T), DLPNO-CCSD(T) and localized ph-AFQMC, for determining the thermochemistry of main group elements. DLPNO-CCSD(T) and localized ph-AFQMC, which offer greater scalability compared to canonical CCSD(T), have emerged over the last decade as pivotal in producing precise benchmark chemical data. Our investigation includes closed-shell, neutral molecules, focusing on their heat of formation and atomization energy sourced from four specific small molecule datasets. Firstly, we selected molecules from the G2 and G3 datasets, noted for their reliable experimental heat of formation data. Additionally, we incorporate molecules from the W4-11 and W4-17 sets, which provide high-level theoretical reference values for atomization energy at 0 K. Our findings reveal that both DLPNO-CCSD(T) and ph-AFQMC methods are capable of achieving a root-mean-square deviation (RMSD) of less than 1 kcal/mol across the combined dataset, aligning with the threshold for chemical accuracy. Moreover, we make efforts to confine the maximum deviations within 2 kcal/mol, a degree of precision that significantly broadens the applicability of these methods in fields such as biology and materials science.
@@
1 1. Introduction
Quantum chemical methods have seen significant improvements in accuracy and computational efficiency when applied to the chemistry of main group elements over the last three decades. Density functional methods, which scale formally with system size $N$ as $N^{4}$ or $N^{3}$ and in practice as $N^{2}$ or even $N$ (for very large systems), can routinely be applied to systems with hundreds to thousands of atoms, with the best functionals providing an average unsigned error on the order of 2–3 kcal/mol for atomization energies of small molecules.1 Alternatively, coupled cluster with perturbative triples (CCSD(T)) has been integrated into numerous quantum chemistry software packages in a robust manner, offering average error rates close to 1 kcal/mol for atomization energies of the same category of small molecules; however, the computational cost scales as $N^{7}$.2, 3 Benchmark methodologies that incorporate higher-order coupled cluster terms and more elaborate treatment of core-valence interactions are capable of producing results reliably within 1 kcal/mol deviation from experimental values. 4 However, the high computational demand of these approaches limits their application to very small molecular systems.
While the accuracy that can be achieved with modern DFT approaches is extremely impressive, DFT calculations on large data sets reveal a significant number of outliers with errors significantly larger than the 2–3 kcal/mol cited above as the average unsigned error. Importantly, outliers can be obscured when large data sets are used and only the average errors are reported.1 In a subsequent paper, we will examine in detail the outlier distribution obtained for a range of modern functionals when compared to benchmark methods and curated experimental results. For the present purposes it is sufficient to note that more work remains to be done to improve the robustness of DFT approaches across a wide range of chemistry, even for main group molecules. Moreover, systems containing transition metals can be prone to a higher incidence of outliers.5
@@
However, in order for wavefunction-based methods to effectively address complex systems, an approach is needed which scales better with system size than the $N^{7}$ of conventional CCSD(T). The past decade has seen the development of two notable methods that address this need, both leveraging the concept of orbital localization—a technique tracing back to Pulay’s work in the 1980s.6 The first is localized coupled cluster (e.g. L-CCSD(T)), the most widely used implementation of which is the DLPNO algorithm of Neese and coworkers.7, 8, 9 The formal scaling of DLPNO-CCSD(T) is $N^{3}$ 10 and extremely impressive timing and accuracy numbers over a wide range of systems (and particularly those restricted to main group chemistry) have been published in the past 5 years.11, 12, 13, 14, 15, 16 DLPNO and related methods (for example other local CCSD(T) methods such as LNO-CCSD(T) 17 and PNO-CCSD(T) 18) represent a revolution in quantum chemical technology as it is now possible to obtain something close to CCSD(T) quality results for systems containing on the order of 100 atoms.
@@
The second approach is auxiliary field quantum Monte Carlo (AFQMC). The AFQMC algorithm was originally developed in the physics community, but it is only in the past 5 years that significant progress has been made in creating a scalable version of the methodology for the ab initio study of molecules. There are several different implementations currently in use. 19, 20, 21, 22. In the present paper, we will focus on two of these implementations. The first is a GPU implementation developed in our groups 23 that exploits localized orbitals 24 in a fashion similar to that employed in L-CCSD(T), reducing the formal scaling of AFQMC from $N^{4}$ to $N^{3}$. We refer to this approach as L-AFQMC. The second 25, 26 is a CPU-based code optimized to enable systematic convergence of the bias in AFQMC energies to near-exact accuracy by using a large number of determinants in the trial function. We designate this implementation as W-AFQMC since it is based on use of the generalized Wick’s theorem. As yet, this method has not been formulated in a localized representation, although work in that direction is ongoing.27 To avoid high computational costs, here we employ W-AFQMC to resolve discrepancies identified in DLPNO-CCSD(T) and L-AFQMC data.
@@
While the computational cost scaling of L-AFQMC and L-CCSD(T) with system size is similar, L-CCSD(T) is considerably faster due to a smaller prefactor. The advantage of AFQMC is that it is formally exact in the limit of the exact trial wavefunction, and in practice, multireference electronic states can often be readily converged due to the ease of utilizing a multiconfigurational trial function.28, 29, 30, 31, 26 This is more crucial for transition metal containing systems than for main group molecules, but there are still main group cases where AFQMC can achieve demonstrably greater accuracy with scalable trial wavefunctions.32, 33 While such trials allow systematic convergence of the bias in phaseless AFQMC (ph-AFQMC), this accuracy comes at the cost of greater computational expense. Designing protocols for generating trials that strike a desired balance between accuracy and cost is an active area of research. Based on our testing with the benchmark sets, we present two approaches with different cost-accuracy trade-offs.
@@
We believe that having two scalable benchmark methodologies with distinct theoretical frameworks offers substantial benefits. These benefits extend not only to the generation of data for evaluating and parameterizing DFT functionals but also to their direct application to challenging systems, like the manganese cluster in Photosystem II. As an example, in one of our recent papers we investigated reactions of organolithium systems relevant to lithium ion batteries.34 We performed DLPNO-CCSD(T) and L-AFQMC calculations to look at both reaction energies and barrier heights. The agreement between the two approaches was remarkably good (a few kcal/mol) across the various reactions that we investigated. We were therefore able to settle on benchmark numbers and use those to evaluate many different DFT functionals, discovering that only a few were able to reproduce the benchmark results reliably. We were then able to use the preferred functionals in computing energies for a large set of organolithium cluster geometries, which we then utilized in parametrizing a machine learning force field (MLFF) for carrying out simulations of lithium ion battery electrolytes.35 While there was no reason, a priori, to doubt the performance of DLPNO-CCSD(T) for these systems, there is very little experimental data available for comparisons, and the validation by a second independent benchmark approach provided a much higher degree of confidence in the results than would otherwise have been possible.
@@
In the present paper, our goal is to provide an assessment of the accuracy of both DLPNO-CCSD(T) and L-AFQMC for main group chemistry atomization energies. We have chosen to focus on atomization energies because (a) a relatively large and reliable data set of benchmark experimental and theoretical values is available for a range of small molecules and (b) atomization energies are one of the most difficult properties for electronic structure methods to compute to high precision, due to the large changes in correlation energy upon atomization, and the relatively minimal cancellation of error. We combine four data sets: the G236, 37 and G338, 39 sets of Pople and coworkers, which claim to have experimental atomization energies that are accurate to 1 kcal/mol or better, and the W4-11 40 and W4-17 41 (TAE - total atomization energy) data sets of Karton, Martin and coworkers, which employ very high level (and hence expensive) theoretical methods to achieve the same level of reliability. The W4-17 set, the latest iteration of the W4 sets, is an extension of the W4-11 set, as is G3 an extension of G2 with the addition of larger molecules. W4-17, however, is restricted to molecules with no more than 8 heavy atoms. On the other hand, the G3 data set contains larger molecules than the W4-17 heavy atom limit and therefore provides a test of how the quantum chemical methods perform for larger systems. There are (as far as we have been able to ascertain) no cases where the errors in the W4-11 and W4-17 results exceed the proposed error bars. In contrast, we have had to update a number of the G2 and G3 experimental values with more recent benchmark values.
@@
The paper is organized as follows. In Section 2, we discuss the data sets including our updating of a number of the reference values in the G2 and G3 sets. We then discuss computational methods including our scalable formulation of AFQMC, DLPNO-CCSD(T), and CCSD(T) methodology (enabling comparisons for the larger systems), basis sets, core-valence corrections, scalar relativistic corrections, extrapolation to the complete basis set limit, and treatment of atomic energies. Section 3 presents the results for all three methods as compared to the relevant experimental or theoretical benchmarks and discusses their implications. In Section 4, we conclude by summarizing our results and discussing future directions. In general, the performance of both DLPNO-CCSD(T) and L-AFQMC are quite robust, with only a few apparent outliers above our targeted 2 kcal/mol threshold. We investigate these outliers using more accurate trials with the W-AFQMC method. This enables us to identify outliers arising from the approximations in the methodology, as opposed to cases that are most likely errors or uncertainties in a few of the experimental reference data. Using this targeted convergence of ph-AFQMC, we are able to produce high-quality atomization energies while minimizing cost over a large dataset.
2 2. Methods
2.1 2.1. Datasets
@@
Table 1: New vs old reference values for heat of formation. The reference values in the column ‘Previous $\Delta_{f}H{(298\,\text{K})}$’ are the same as used in the original G2/G3 test set papers.37, 39 All values are reported in kcal/mol. Where the source reports an experiment uncertainty, we have included the uncertainty in the table along with the value.
Reference values for the W4 sets are high-level theoretical atomization energies at 0 K, excluding zero-point energy (ZPE) (labeled as the property TAEe by Karton et al. 41). The reference values for the G2 and G3 sets are experimental heats of formation at 298 K, and the same as those used by Curtiss et al. 36, 38. However, there are some exceptions where we have found conflicting values, as summarized in Table 1, where each of the sources of the updated reference is listed. The reported reference values from ATcT42, 43 postdate the G2 and G3 papers. Furthermore, ATcT collates the most recent experiments and theory from various sources using a self-consistent approach and numerous different reactions44 and is readily updated 45. Therefore, where ATcT data is available and conflicting with the reference used by Curtiss et al., we instead use the ATcT value. In addition, for the cases of AlCl3 and AlF3, we found disagreement between the experimental heats of formation and the W4 reported atomization energy. Moreover, the heat of formation values of these molecules are not present in the ATcT database. Thus, for the updated reference values of these molecules, we have converted the W4 atomization values into heats of formation at 298 K using temperature corrections from DFT (refer to Section 2.9 for details). For the three cases of LiF, azulene, and benzoquinone, where no ATcT or W4 value is available, we use the references reported by the NIST database CCCBDB46 instead of those reported by Curtiss. As shown in Table 1, the two sources also give conflicting heats of formation for these three molecules. All of our benchmark wavefunction methods yield results that are within 2 kcal/mol of the latter source, rather than the former reference values. Although this choice is not based on information about the reference alone, CCSD(T) and AFQMC offer independent evaluations of the experimental data, especially in the case of conflict. In particular, we converge AFQMC with respect to the number of determinants for these cases using W-AFQMC. We refer the reader to Section 2.9 for the method of obtaining deviations against reference values.
2.3 2.3. Phaseless AFQMC Formulation
@@
$h_{pq}$ are one-electron integrals and $V_{pqrs}=\left(pr|qs\right)=\langle pq|rs\rangle{}{}$ (chemists’ and physicists’ notation respectively) are two-electron integrals. Numerically, we propagate
@@
where $p(\mathbf{x})$ is a Gaussian probability density function and $\hat{B}(\mathbf{x})$ is a one-body propagator depending on the auxiliary fields $\mathbf{x}$. This multidimensional integral is evaluated using Monte Carlo importance sampling to obtain a stochastic representation of the wave function. For a more in-depth description of AFQMC, we refer the reader to these review articles 47, 48.
@@
Due to the fermionic sign problem, the signal-to-noise ratio generally decays exponentially during the imaginary time propagation. It is possible to eliminate the sign problem using a constraint at the expense of a bias in the resulting energies. In this work, we use a constraint referred to as the phaseless approximation (ph-AFQMC), where the phase of the walkers is restricted according to a trial wavefunction. The bias induced by the trial wavefunction can be systematically reduced by improving its quality, for example, by increasing the active space or number of determinants included in the trial. The bias is formally zero in the limit of the exact trial (see the next subsection, Section 3.3, and SI Section 3 for trial details). We refer the reader to our previous work24 for our approach to localized ph-AFQMC (L-AFQMC) which involves compressing the electron repulsion integrals in the localized orbital basis. Effectively, the scaling for the energy evaluation, the steepest scaling step with system size, is $N^{2}M+N_{\text{det}}N^{2}$ (with a prefactor depending on the compressed electron repulsion integral tensor), where $M$ is the number of basis functions, $N$ is the number of electrons, and $N_{\text{det}}$ is the number of determinants. See SI Section 4 for an estimate of localization error. For the practical deployment of L-AFQMC, we have developed two protocols, AFQMC 0 and AFQMC I, which are discussed in detail below as well as in Section 3.3. AFQMC I is a scalable AFQMC protocol (scaling $\sim N^{3}$) that achieves an accuracy comparable to that of DLPNO-CCSD(T), albeit with a significantly larger prefactor for the computation time. AFQMC 0 uses a black-box but less elaborate trial function, but is less accurate, particularly for molecules with significant multireference character.
@@
We also present results for a selected subset of molecules computed with another implementation of AFQMC 25, 26. This method, which uses a generalized Wick’s theorem approach to efficiently evaluate energies with mutlideterminantal trials, will be referred to in what follows as W-AFQMC. The advantage of this approach is that it scales as $MN_{\text{det}}+N^{2}M^{2}$, which allows the use of a larger number of determinants in the trial wavefunction at an accessible computational cost. It enables one to converge the phaseless bias to the near-exact limit in a given basis set by increasing the number of active space orbitals and determinants included in trial wavefunctions. We use W-AFQMC to calculate energies of the outliers obtained from AFQMC I using on the order of 10,000 determinants and refer to the results in which the L-AFQMC outliers are replaced by the W-AFQMC results as AFQMC II (along with a select few other cases, see SI Section 3). This approach helps us to more confidently address the question, discussed in detail below, as to whether the AFQMC I outliers are due to the phaseless bias or more likely the result of errors in the reference data. We also run apparent DLPNO-CCSD(T) and CCSD(T) outliers with AFQMC II to assess their status.
2.4 2.4. AFQMC Trial Generation
@@
In this work, we use a procedure to generate multi-determinant HCI (heat bath configuration interaction) and HCISCF trials for the entire dataset. The CAS family of trials (in this work, we use the HCI solver49, 50) provides a robust way of including static correlation in the reference of AFQMC and has been shown19, 28 to generally perform more accurately than single determinant trials. Nonetheless, the selection of active space for these methods is non-trivial. Akin to multireference perturbation theory methods, a common practice is to use an active space, usually minimal, based on chemical intuition and to pick the leading determinants from the expansion in this active space. However, this does not systematically provide chemically accurate energies. Here, we generally follow a two-step process to select the active space in a relatively automated way that can be applied to large datasets. First, HCI is performed on a “valence” active space, selected based on the atomic composition of the molecule. See SI Section 3 and main text Section 3.3 for the considerations for selecting this active space. Using the spatial 1-RDM of the resulting state, we calculate the natural orbitals and their occupation numbers (NOONs) in the natural orbital basis. We choose a subset of active orbitals from this set based on a NOON threshold ($\delta$) as $\delta\leq\text{NOON}\leq 2-\delta$. This procedure is often used to flag the more correlated group of orbitals in quantum calculations. A second HCI calculation (or HCISCF for W-AFQMC trials, refer to the SI Section 3) is then performed with the second smaller active space, and this forms the final trial wavefunction. Generally, we choose the number of determinants necessary to retain 99.5% of the CI weight from this final trial wavefunction unless indicated otherwise. Briefly, AFQMC 0 is fully automated and uses L-AFQMC, and thresholds are chosen to be loose, and for the outliers with AFQMC 0, AFQMC I combines trials with larger active spaces, and AFQMC II in turn combines trials with even larger active spaces run with W-AFQMC. This progression gives some indication of the number of determinants required to converge each molecule, but not much more than necessary. For more details, refer to the discussion in Section 3 and SI Section 3.3.
@@
All L-AFQMC trials were generated with the PySCF package51 where we obtain the Hamiltonian and electron repulsion integrals. The HCI trials are obtained using Dice49, 50, 52 in conjunction with PYSCF. L-AFQMC energy is measured in blocks of size 0.1 Ha-1 of 20 timesteps each (each timestep being 0.005 Ha-1). In total, we propagate for between 2000 and 3000 such blocks for each molecule, with 1920 total walkers. W-AFQMC calculations used the same time-step of 0.005 Ha-1 for molecules containing only first row atoms, and used a time step of 0.0025 Ha-1 for those with heavier atoms. For W-AFQMC, we use 250 walkers and propagate for 1000 blocks of 50 timesteps each.
2.5 2.5. CCSD(T) and DLPNO-CCSD(T)